Three blocks A,B,C of masses 4,2,1kg respectively are in contact on a frictionless surface. If a force of 14N is applied first block A, then the contact force between A and B is what? 

Grade:12

1 Answers

Arun
5 years ago
 
the acceleration due to force 14N=
 
F=ma
F/m =a
14/(4+2+1)
14/7
=2m/s2
 
the contact force between A and B is used to move the bodies B and C only.
F=ma
F=(2+1)2
F=3×2
=6N